# Permissioned Blockchain in Advanced Air Mobility: A Performance Analisys for UTM

### v1

@@ Abstract
The rapid adoption of Uncrewed Aerial Vehicles (UAVs) has driven aviation authorities to propose distributed Uncrewed Traffic Management (UTM) architectures. Several studies have advocated blockchain as a promising technology to meet these requirements. However, since UTM is a safety-critical and highly regulated domain, compliance with standards and regulatory frameworks is as crucial as performance and security. This work benchmarks two distributed architectures aligned with current regulatory frameworks: the Linux Foundation’s InterUSS platform and a Hyperledger Fabric-based private ledger. Our findings reveal that blockchain-based systems require architectures specifically designed for aeronautical performance constraints.
@@ VI Conclusion and Future Works
This work compared two distributed architectures aligned with current UTM frameworks: the federated InterUSS Platform and a permissioned blockchain based on Hyperledger Fabric. By benchmarking the Operational Intent Reference (OIR) registration process, the study showed that both systems can handle moderate workloads, but their scalability diverges beyond 30 TPS. InterUSS sustained sub-second latency and stable tail behavior, while Fabric exhibited exponential degradation due to its endorsement and ordering processes. These results indicate that blockchain-based systems, although valuable for ensuring immutability and accountability, still face challenges in meeting the timing constraints required for aeronautical-grade operations.
Future work will explore the use of blockchain oracles integrated with Hyperledger Besu to improve interoperability and real-time responsiveness. Since Besu implements the Ethereum Virtual Machine (EVM), it allows compatibility with widely used development frameworks and smart-contract ecosystems. This approach may enable hybrid architectures that combine the auditability and trust of DLT with the agility and scalability of federated UTM environments.

### v2

# Permissioned Blockchain in Advanced Air Mobility: A Performance Analysis for UTM

@@ Abstract
The integration of Uncrewed Aerial Vehicles (UAVs) into low-altitude airspace has led authorities to adopt distributed Uncrewed Traffic Management (UTM) architectures that ensure interoperability and safety. Blockchain has been proposed as an enabler for trustworthy coordination among UTM stakeholders. Yet, its real-time performance under aeronautical constraints remains insufficiently characterized. This paper presents a quantitative benchmark comparing two regulation-compliant distributed architectures: the federated InterUSS platform maintained by the Linux Foundation and a permissioned blockchain based on Hyperledger Fabric. Both systems were evaluated through Operational Intent Reference (OIR) registration workloads generated via Hyperledger Caliper, measuring throughput, latency, and transaction loss under loads up to 50 transactions per second. Results show that InterUSS sustained sub-second latency and stable performance up to 30 TPS. At the same time, Fabric exhibited exponential degradation with median latency exceeding 3 s and tail latencies above 15 s beyond that point. These findings demonstrate that blockchain-based architectures must be redesigned to meet aeronautical timing and scalability requirements, suggesting that hybrid models combining distributed ledgers for auditability with federated frameworks for real-time coordination are more suitable for future UTM deployments.
@@ VI Conclusion and Future Works
This work analyzed smart city drones solutions: the federated InterUSS platform and a permissioned blockchain based on Hyperledger Fabric. InterUSS sustained sub-second latency up to 30 TPS with stable tails, while Fabric exceeded 3 s median latency and 15 s p90 beyond this load, saturating near 50 TPS. These results show that current permissioned-ledger designs cannot meet aeronautical real-time constraints. A hybrid model that couples federated coordination with DLT-based auditability emerges as a more practical path for scalable, regulation-compliant UTM backends. In future work we plan extend the benchmark to mixed workloads and larger validator sets, and explore hybrid designs using blockchain oracles and Besu-based smart contracts in UTM.